# Boundary layers in thermal convection are fluctuation-dominated

### v1

@@ III Mean temperature and velocity profiles
for $i=x,y$. Figure 3 displays the result of this analysis for $Ra=10^{9}$. In panel (a), the mean profile of the $x$-velocity component is shown as a function of the averaging time which was varied from $\Delta t=400\,T_{f}$ (= 20 snapshots) to $1600\,T_{f}$ (= 80 snapshots). The profile converges steadily to zero, though not uniformly. A fit of the mean profiles to the two-dimensional Blasius solution $f(\eta)$ has been applied subsequently; the solution $f(\eta)=\psi(x,z)/\sqrt{xU_{\infty}\nu}$ and $\eta=z/\delta(x)$ (Schlichting and Gersten, 2016). Here, $\psi(x,z)$ is the stream function and $u_{x}=\partial\psi/\partial z$ and $u_{z}=-\partial\psi/\partial x$. Recall that at distance $\eta=5$ the Blasius profile reaches a streamwise velocity magnitude of $f^{\prime}(\eta)=0.99U_{\infty}$.
In the absence of a definable leading edge distance $x$, we match $\langle u_{x}\rangle_{A,t}(z)/U_{\infty}$ and $z/\delta_{\infty}$ to $f^{\prime}(\eta)$ which is reported in panels (b,c) of figure 3. Therefore, the numerical profiles are rescaled such that the first local maximum of $\langle u_{x}\rangle_{A,t}$ corresponds to $U_{\infty}$. The fits are shown again for different time intervals. It is seen in panel (c) that the velocity scale $U_{\infty}$ drops to a very small magnitude $\sim 10^{-3}U_{f}$ while $\delta_{\infty}\lesssim 0.1H$ in case of $Ra=10^{9}$. In table 2, the results for all Rayleigh numbers and both horizontal components are listed. No clear trend of the velocity $U^{x,y}_{\infty}$ with Rayleigh number is detectable, the magnitude is between ${\cal O}(10^{-3})$ and ${\cal O}(10^{-2})$. The boundary layer thickness parameters vary as well when $x$- and $y$-directions are compared at fixed $Ra$. They decrease with increasing Rayleigh number. Furthermore, we calculate the corresponding shear Reynolds numbers $Re_{\rm shear}=U_{\infty}\delta_{\infty}/\nu$, which are found to be very small for all cases. We have also repeated this analysis for a combination of both components (not shown), though not for the magnitude which would become insensitive to different directions of local flow patterns. Panel (d) of figure 3 shows the mean profiles for all 7 simulation runs.
Panels (e) and (f) show the vertical profiles of root-mean-square velocities in red,
@@
together with profiles obtained in the individual snapshots in gray.
@@
Table 2: Fit parameters $U^{x}_{\infty}$ and $U^{y}_{\infty}$ as well as the corresponding thicknesses $\delta^{x}_{\infty}$ and $\delta^{y}_{\infty}$ of the combined plane-time averaged mean profiles of the horizontal velocity components $\langle u_{x}\rangle_{A,t}$ and $\langle u_{y}\rangle_{A,t}$ to the Blasius profile $U_{\infty}f^{\prime}(\eta)$ with $\eta=z/\delta_{\infty}$. The corresponding shear Reynolds numbers $Re^{x,y}_{\mathrm{shear}}=U^{x,y}_{\infty}\delta^{x,y}_{\infty}/\nu$ are also given.
@@
Figure 4 summarizes mean profiles for all 7 simulation runs, the mean temperature profile and the root-mean-square profiles of temperature, horizontal velocity components, and all three velocity components. The temperature fluctuation profile is obtained by $T_{\rm rms}(z)=\langle\theta^{2}\rangle^{1/2}_{A,t}$ with $\theta({\bm{x}},t)=T({\bm{x}},t)-\langle T\rangle_{A,t}(z)$. The correspondingly related characteristic scales are indicated by horizontal dashed lines and detailed in table 3. It is seen that the thermal boundary layer thickness $\delta_{T}=1/(2Nu)$ is slightly larger than the distance from the wall of the maximum of the temperature fluctuation profile, $\delta_{T,{\rm rms}}$. Increasingly larger are distances from the wall to the maxima of the velocity fluctuation profiles, obeying a ratio of $\delta_{U,{\rm rms}}/\delta_{T,{\rm rms}}\approx 2$ for $Ra=10^{5}$ up to approximately 14 for $Ra=10^{11}$; see panel (d).
@@
Figure 4: Mean temperature profiles compared to each other for all 7 runs. We show the mean profile of temperature $\langle T\rangle_{A,t}(z)$ in panel (a), root-mean-square profile of temperature $T_{\rm rms}(z)$ in (b), root-mean-square profile with respect to the horizontal velocity components, $U^{h}_{\rm rms}(z)$ in (c), and with respect to all three components $U_{\rm rms}(z)$ in (d). The corresponding height scales of these profiles are indicated by dashed horizontal lines and listed in table 3.
$Ra$ $\delta_{T}$ $\delta_{T,{\rm rms}}$ $\delta^{h}_{U,{\rm rms}}$ $\delta_{U,{\rm rms}}$
$10^{5}$ ($1.17\pm 0.04)\times 10^{-1}$ ($1.27\pm 0.08)\times 10^{-1}$ ($1.23\pm 0.07)\times 10^{-1}$ ($2.5\pm 1.4)\times 10^{-1}$
$10^{6}$ ($6.13\pm 0.15)\times 10^{-2}$ ($6.13\pm 0.27)\times 10^{-2}$ ($8.80\pm 0.54)\times 10^{-2}$ ($1.8\pm 1.2)\times 10^{-1}$
$10^{7}$ ($3.21\pm 0.06)\times 10^{-2}$ ($3.07\pm 0.10)\times 10^{-2}$ ($6.99\pm 0.58)\times 10^{-2}$ ($1.07\pm 0.51)\times 10^{-1}$
$10^{8}$ ($1.64\pm 0.02)\times 10^{-2}$ ($1.52\pm 0.04)\times 10^{-2}$ ($5.08\pm 0.35)\times 10^{-2}$ ($6.51\pm 0.92)\times 10^{-2}$
$10^{9}$ ($8.22\pm 0.12)\times 10^{-3}$ ($7.47\pm 0.18)\times 10^{-3}$ ($3.52\pm 0.28)\times 10^{-2}$ ($4.3\pm 2.0)\times 10^{-2}$
$10^{10}$ ($4.06\pm 0.05)\times 10^{-3}$ ($3.72\pm 0.17)\times 10^{-3}$ ($2.32\pm 0.20)\times 10^{-2}$ ($2.82\pm 0.37)\times 10^{-2}$
$10^{11}$ ($1.96\pm 0.02)\times 10^{-3}$ ($1.77\pm 0.05)\times 10^{-3}$ ($2.12\pm 0.24)\times 10^{-2}$ ($2.61\pm 0.37)\times 10^{-2}$
@@
Table 3: Differently defined boundary layer heights of temperature and velocity according to the vertical mean profiles plotted in figure 4. The standard deviation accompanies each mean value.
IV Scaling laws of fluctuations with Rayleigh number
@@
Figure 5: Scaling of the root-mean-square velocities and temperature with respect to Rayleigh number. All values are obtained by a combined average with respect to volume and time. (a) Horizontal velocity components. (b) Vertical velocity component. (c) All three velocity components. (d) Temperature. Dashed lines denote power law fits to the data. Blue open circles stand for data obtained in the bulk volume $V_{b}$. In panels (b–d), we add data from Iyer et al. (2020) with separate fits. The corresponding temperature data are obtained for the cylindrical cell interior, $r\leq 0.03H$ and $Ra\geq 10^{9}$.
@@
Figure 5 summarizes the root-mean-square fluctuations of the three velocity components and the temperature. They are obtained by a combined average with respect to the full volume $V=L^{2}H$ and time, e.g. $u_{x,{\rm rms}}=\langle u_{x}^{2}\rangle^{1/2}_{V,t}$. The quantity $U_{\rm rms}$ denotes again the fluctuations with respect to all three velocity components. It is seen that the dependence of the velocity fluctuations on the Rayleigh number is very weak with $\beta\lesssim 0.042$. The temperature fluctuations drop with a smaller power law exponent, $T_{\rm rms}\sim Ra^{-\beta}$, which is found to be $\beta=0.118$ for the present data. This exponent is slightly smaller in magnitude than those reported in experiments in cylindrical cells of aspect ratio 1/2. For comparison, Castaing et al. (1989), Niemela et al. (2000) and Wu and Libchaber (1992) report exponents of $\beta\approx 0.145$. We also analysed the temperature fluctuations in the bulk of the layer, which takes a volume average with respect to $V_{b}=L^{2}\times[0.4,0.6]$ and time. The exponent changes to $\beta=0.140$ which is closer to the experiments. We have verified that a variation of the thickness of the bulk volume $V_{b}$ does not alter the results significantly.
@@
In panels (b–d) of figure 5, we added data from the DNS of Iyer et al. (2020) for comparison, which were obtained in a closed slender cylindrical cell of aspect ratio $\Gamma=0.1$. It is seen that exponents of the power law fits are close to those of the present simulation series. The prefactors differ as expected, because the former DNS data were obtained for a geometrically constrained convection flow. This finding of nearly the same scaling exponents suggests a robust geometry-independent trend of all fluctuations with respect to the Rayleigh number. Geometry-specific aspects mostly affect the prefactor.
V Decomposition into coherent and incoherent boundary layer regions
@@
The orientation of the boundary layer flow varies strongly as shown in figures 6 (a) and (d), where we plot the pointwise orientation angle of the horizontal velocity $\varphi=\arctan(u_{x}/u_{y})\in[-\pi,\pi]$ for a snapshot at $Ra=10^{10}$ at $z=\delta_{T}$ and $z=H-\delta_{T}$, respectively. At both these heights, we break down the horizontal plane into a cover of $10^{4}$ disjoint square boxes of area content $A_{i}=A/10^{4}$, where $A=L^{2}$. We then calculate the mean horizontal velocity $\bar{\bm{u}}_{h}(A_{i})$ in each of the $A_{i}$ and decompose the cross section into coherent and incoherent boundary layer regions for $|\bar{\bm{u}}_{h}(A_{i})|>U^{h}_{\rm rms}(\delta_{T})$ and $|\bar{\bm{u}}_{h}(A_{i})|\leq U^{h}_{\rm rms}(\delta_{T})$, respectively. Panels (b, e) of figure 6 show that coherent shear-dominated patches separated by incoherent flow regions (in gray). The superposed streamlines indicate the different flow orientations of the shear-dominated regions. Panels (c,f) show the corresponding snapshots of the temperature field $T$ at $z=0.1$ above the bottom and $z=0.9$ below the top which is a distance of 25 $\delta_{T}$ away from the walls.
It is clearly seen that the hotter regions at $z=0.1$ and the colder regions at $z=0.9$ coincide with the incoherent flow regions at the edge of the thermal boundary layer. This is the imprint of the turbulent superstructure of convection in this setup at $\Gamma=4$. We have also determined that the area fraction of the incoherent regions remains nearly constant at approximately 60% of $A$ for the whole Rayleigh number range, which further supports its connection to the turbulent superstructure that consists of a few large-scale circulation rolls which extend over the periodic horizontal boundaries. The physical interpretation is that the incoherent regions correspond to the upwelling (downwelling) motion of clustered plumes outside shear dominated patches. It suggests a vertical coarsening from the plate into the bulk, which remains to be analysed in the future.
@@
Figure 6: Boundary layer flow structure for $Ra=10^{10}$ at the top and bottom walls. (a) Snapshot of the pointwise orientation angle of the horizontal velocity vector at $z=\delta_{T}$. (b) Decomposition of the same snapshot into coherent and incoherent (in gray) flow regions at $z=\delta_{T}$. For the coherent shear-dominated regions, we overlay horizontal streamlines and indicate their local flow direction. (c) Corresponding temperature field at $z=0.1\approx 25\delta_{T}$. (d) Same snapshot with the pointwise orientation angle of the horizontal velocity vector at $z=H-\delta_{T}$. (e) Decomposition similar to (b) for $z=H-\delta_{T}$. (f) Corresponding temperature field at $z=0.9\approx 25\delta_{T}$.
VI Conclusions

### v2

@@ 3 Mean profiles of temperature and velocity
for $i=x,y$ and $N_{t}$ the number of snapshots. For all runs, $N_{t}=20$ to obtain equidistant and statistically independent realizations of the flow. Figure 4 displays the result of this analysis for $Ra=10^{9}$. In panel (a), the mean profile of the $x$-velocity component is shown as a function of even longer averaging times, which were varied from $\tau_{\rm total}=400\,T_{f}$ (= 20 snapshots) to $1600\,T_{f}$ (= 80 snapshots). The profile converges steadily to zero, though not uniformly. There is essentially no mean flow. If we insist upon fitting the near wall mean profiles to the two-dimensional Blasius solution, for example, the result is shown in panel (b). In the absence of a definable leading edge distance $x$, we match $\langle u_{x}\rangle_{A,t}(z)/U_{\infty}$ and $z/\delta_{\infty}$ to $f^{\prime}(\eta)$ which is reported in panels (b,c) of figure 4. Recall that the Blasius solution $f(\eta)=\psi(x,z)/\sqrt{xU_{\infty}\nu}$ and $\eta=z/\delta(x)$ (Schlichting & Gersten, 2016), where $\psi(x,z)$ is the stream function and $u_{x}=\partial\psi/\partial z$ and $u_{z}=-\partial\psi/\partial x$. The numerical profiles are rescaled such that the first local maximum of $\langle u_{x}\rangle_{A,t}$ corresponds to $U_{\infty}$. The fits are shown for different time intervals in (b). Recall that at distance $\eta=5$ the Blasius profile reaches a streamwise velocity magnitude of $f^{\prime}(\eta)=0.99U_{\infty}$. In this case, panel (a) shows that the maximum velocity reached is of the order 1% or less of the free-fall velocity, which is the only characteristic velocity of the flow. As shown in panel (c), no clear trend of the velocity $U^{x,y}_{\infty}$ with averaging time is detectable, and the magnitude is between ${\cal O}(10^{-3})$ and ${\cal O}(10^{-2})$. In table 2, the results for all Rayleigh numbers and both horizontal components are listed.
The boundary layer thickness parameters vary when $x$- and $y$-directions are compared at fixed $Ra$. They decrease with increasing Rayleigh number. Furthermore, we calculate the corresponding shear Reynolds numbers $Re_{\rm shear}=U_{\infty}\delta_{\infty}/\nu$, which are found to be very small for all cases. Panels (d) and (e) of figure 4 show the mean profiles for all 7 simulation runs and for both horizontal velocity components. They underline the very small mean-flow amplitudes for all Rayleigh numbers in this series. Panel (f) of the figure compares three runs at $Ra=10^{9}$, which are indicated with a dagger symbol in table 1, at aspect ratios $\Gamma=2$, 4, and 8. Again, the mean flow amplitudes $\langle u_{x}\rangle_{A,t}(z)$ are comparable and very small such that an aspect ratio dependence for this result can be excluded when periodic boundary conditions in the horizontal directions are applied. The non-monotonic behavior is simply a reflection of the long averaging times required for convergence in convection studies.
4 Fluctuation profiles of temperature and velocity
The mean vertical profiles of the root-mean-square velocities are given by
@@ 4 Fluctuation profiles of temperature and velocity
where we distinguish between horizontal and full profiles. The fluctuation profiles for the velocity fields are obtained from the full components since the means are so small. We have verified that the differences in the procedure produce only very small changes.
@@
[…]
@@
Figure 5 summarizes mean profiles for 7 simulation runs, the mean temperature profile and the root-mean-square profiles of temperature, horizontal velocity components, and all three velocity components. The temperature fluctuation profile is similarly obtained by $T_{\rm rms}(z)=\langle\theta^{2}\rangle^{1/2}_{A,t}$ with $\theta({\bm{x}},t)=T({\bm{x}},t)-\langle T\rangle_{A,t}(z)$. The correspondingly related characteristic scales are indicated by horizontal dashed lines and detailed in table 3. It is seen that the thermal boundary layer thickness $\delta_{T}=1/(2Nu)$ is slightly larger than the distance from the wall of the maximum of the temperature fluctuation profile, which we term the thermal fluctuation boundary layer thickness or, for short, thermal fluctuation thickness, $\delta_{T,{\rm rms}}$. Increasingly larger are distances from the wall to the maxima of the velocity fluctuation profiles, obeying a ratio of $\delta_{U,{\rm rms}}/\delta_{T,{\rm rms}}\approx 2$ for $Ra=10^{5}$ up to approximately 13 for $Ra=10^{11}$; see panel (d). The corresponding thicknesses are termed velocity fluctuation thickness.
Furthermore, we repeated the fluctuation profile analysis for the dependence on different aspect ratios, one on each side of 4, at $Ra=10^{9}$. These runs are indicated again by a dagger symbol in table 3, where we collect the corresponding thickness scales. The corresponding profiles are shown in figure 6. While the temperature profiles collapse close to the walls, thus displaying no sensitivity with respect to the aspect ratio $\Gamma$ in this range, the velocity profiles are affected by $\Gamma$. However, the resulting velocity fluctuation thicknesses are found to agree well for $\Gamma=4$ and 8 (by 1.5% or less), in terms of both horizontal and full velocity fluctuations. The finding supports our considered view that $\Gamma\geq 4$ is sufficient to obtain horizontal homogeneity for the statistics already introduced.
@@
Figure 5: Mean profiles are compared to each other for all 7 runs. We show the mean profile of temperature $\langle T\rangle_{A,t}(z)$ in panel (a), root-mean-square profile of temperature $T_{\rm rms}(z)$ in (b), root-mean-square profile with respect to the two horizontal velocity components, $U^{h}_{\rm rms}(z)$, in (c), and with respect to all three components, $U_{\rm rms}(z)$, in (d). The corresponding height scales of these profiles are indicated by dashed horizontal lines in the figures and listed in table 3. Note that for the two lowest $Ra$, they exceed the displayed plot range of the figure in panel (d).
Figure 6: Mean temperature profiles compared to each other for different aspect ratio at $Ra=10^{9}$. Similar to figure 5, we show the mean profile of temperature $\langle T\rangle_{A,t}(z)$ in panel (a), root-mean-square profile of temperature $T_{\rm rms}(z)$ in (b), root-mean-square profile with respect to the horizontal velocity components, $U^{h}_{\rm rms}(z)$ in (c), and with respect to all three components $U_{\rm rms}(z)$ in (d). The corresponding height scales of these profiles are indicated by dashed horizontal lines and are also listed in table 3.
$Ra$ $\Gamma$ $\delta_{T}$ $\delta_{T,{\rm rms}}$ $\delta^{h}_{U,{\rm rms}}$ $\delta_{U,{\rm rms}}$
$10^{5}$ 4 $(1.17\pm 0.04)\times 10^{-1}$ $(1.27\pm 0.08)\times 10^{-1}$ $(1.23\pm 0.07)\times 10^{-1}$ $(2.39\pm 1.32)\times 10^{-1}$
$10^{6}$ 4 $(6.13\pm 0.15)\times 10^{-2}$ $(6.13\pm 0.27)\times 10^{-2}$ $(8.80\pm 0.54)\times 10^{-2}$ $(1.74\pm 1.12)\times 10^{-1}$
$10^{7}$ 4 $(3.21\pm 0.06)\times 10^{-2}$ $(3.07\pm 0.10)\times 10^{-2}$ $(6.99\pm 0.58)\times 10^{-2}$ $(1.07\pm 0.51)\times 10^{-1}$
$10^{8}$ 4 $(1.64\pm 0.02)\times 10^{-2}$ $(1.52\pm 0.04)\times 10^{-2}$ $(5.08\pm 0.35)\times 10^{-2}$ $(6.51\pm 0.92)\times 10^{-2}$
$10^{9}$ 4 $(8.22\pm 0.12)\times 10^{-3}$ $(7.47\pm 0.18)\times 10^{-3}$ $(3.52\pm 0.28)\times 10^{-2}$ $(4.33\pm 1.97)\times 10^{-2}$
$10^{9\,\dagger}$ 2 $(8.01\pm 0.15)\times 10^{-3}$ $(7.07\pm 0.29)\times 10^{-3}$ $(2.94\pm 0.49)\times 10^{-2}$ $(3.75\pm 0.72)\times 10^{-2}$
$10^{9\,\dagger}$ 4 $(8.27\pm 0.10)\times 10^{-3}$ $(7.44\pm 0.21)\times 10^{-3}$ $(3.28\pm 0.28)\times 10^{-2}$ $(4.15\pm 0.63)\times 10^{-2}$
$10^{9\,\dagger}$ 8 $(8.20\pm 0.04)\times 10^{-3}$ $(7.51\pm 0.12)\times 10^{-3}$ $(3.23\pm 0.12)\times 10^{-2}$ $(4.20\pm 0.22)\times 10^{-2}$
$10^{10}$ 4 $(4.06\pm 0.04)\times 10^{-3}$ $(3.67\pm 0.11)\times 10^{-3}$ $(2.39\pm 0.19)\times 10^{-2}$ $(2.79\pm 0.27)\times 10^{-2}$
$10^{11}$ 4 $(1.97\pm 0.02)\times 10^{-3}$ $(1.81\pm 0.05)\times 10^{-3}$ $(2.07\pm 0.27)\times 10^{-2}$ $(2.55\pm 0.39)\times 10^{-2}$
@@
Table 3: Differently defined boundary layer thicknesses of temperature and velocity according to the vertical mean profiles plotted in figure 5. The standard deviation accompanies each mean value.
5 Scaling of combined volume-time averaged fluctuation with Rayleigh number
@@
Figure 7: Scaling of the root-mean-square velocities and temperature with respect to Rayleigh number. All values are obtained by a combined average with respect to volume and time. (a) Horizontal velocity components. (b) Vertical velocity component. (c) All three velocity components. (d) Temperature. Dashed lines denote power law fits to the data. Blue open circles stand for data obtained in the bulk volume $V_{b}$. In panels (b–d), we add data from Iyer et al. (2020) with separate fits. The corresponding temperature data are obtained there for the cylindrical cell interior, $r\leq 0.03H$ and $Ra\geq 10^{9}$.
@@
Figure 7 summarizes the root-mean-square fluctuations of the three velocity components and the temperature. They are obtained by a combined average with respect to the full volume $V=L^{2}H$ and time, e.g. $u_{x,{\rm rms}}=\langle u_{x}^{2}\rangle^{1/2}_{V,t}$. The quantity $U_{\rm rms}$ denotes again the fluctuations with respect to all three velocity components. It is seen that the dependence of the velocity fluctuations on the Rayleigh number is very weak with $\beta\lesssim 0.042$. The temperature fluctuations drop with a smaller power law exponent, $T_{\rm rms}\sim Ra^{-\beta}$, which is found to be $\beta=0.119$ for the present data. This exponent is slightly smaller in magnitude than those reported in experiments in cylindrical cells of aspect ratio 1/2. For comparison, Castaing et al. (1989), Niemela et al. (2000) and Wu & Libchaber (1992) report exponents of $\beta\approx 0.145$. We also analysed the temperature fluctuations in the bulk of the layer, which takes a volume average with respect to $V_{b}=L^{2}\times[0.4,0.6]$ and time. The exponent changes to $\beta=0.141$ which is closer to the experiments. We have verified that a variation of the thickness of the bulk volume $V_{b}$ does not alter the results significantly.
@@
In panels (b–d) of figure 7, we added data from the DNS of Iyer et al. (2020) for comparison, which were obtained in a closed slender cylindrical cell of aspect ratio $\Gamma=0.1$. It is seen that exponents of the power law fits are close to those of the present simulation series. The prefactors differ as expected, because the former DNS data were obtained for geometrically constrained convection. This finding of nearly the same scaling exponents suggests a robust geometry-independent trend of all thermal fluctuations with respect to the Rayleigh number. Geometry-specific aspects mostly affect the prefactor.
6 Decomposition into coherent and incoherent boundary layer regions
@@
The orientation of the boundary layer flow varies strongly as shown in figures 8 (a) and (c), where we plot the orientation angle of the horizontal velocity $\varphi=\arctan(u_{y}/u_{x})\in[-\pi,\pi]$ for a snapshot at $Ra=10^{10}$ at $z=\delta_{T}$ and $z=H-\delta_{T}$, respectively. At both these heights, we cover the horizontal plane into $10^{4}$ disjoint square boxes of area content $A_{i}=A/10^{4}$, where $A=L^{2}$. We then calculate the mean horizontal velocity $\bar{\bm{u}}_{h}(A_{i})$ in each of the $A_{i}$ and decompose the cross section into coherent and incoherent boundary layer regions for $|\bar{\bm{u}}_{h}(A_{i})|>U^{h}_{\rm rms}(\delta_{T})$ and $|\bar{\bm{u}}_{h}(A_{i})|\leq U^{h}_{\rm rms}(\delta_{T})$, respectively. Similar decompositions have been applied to analyse the spatio-temporal intermittency of the transition to turbulence in shear flow turbulence in extended domains, see e.g. Hof (2022). Panels (a, c) of figure 8 show that coherent shear-dominated patches are separated by incoherent flow regions (in gray). See also figure 1. The superposed streamlines indicate the different flow orientations of the shear-dominated regions.
Panels (b,d) of the same figure show the corresponding snapshots of the temperature field $T$ at $z=0.1$ above the bottom and $z=0.9$ below the top, which are distances of 25 $\delta_{T}$ away from the walls at $Ra=10^{10}$. It is clearly seen that the hotter regions at $z=0.1$ and the colder regions at $z=0.9$, both of which are displayed in gray, coincide fairly well with the incoherent flow regions at the edge of the thermal boundary layer. We can define overlap factors $0\leq\tilde{O}\leq 1$ by
$\tilde{O}_{\rm bot}=\frac{A_{\rm incoh}\bigcup A_{T>T_{0}}}{\max(A_{\rm incoh},A_{T>T_{0}})}\quad\mbox{and}\quad\tilde{O}_{\rm top}=\frac{A_{\rm incoh}\bigcup A_{T<T_{0}}}{\max(A_{\rm incoh},A_{T<T_{0}})}\,,$ (7)
with $T_{0}=0.5$. Here, we find mean overlaps of $\langle\tilde{O}_{\rm bot}\rangle=0.60$ and $\langle\tilde{O}_{\rm top}\rangle=0.63$, where the average is taken over the snapshots. The physical interpretation is as follows: the incoherent regions correspond to dominant hotter upwelling (colder downwelling) motions. These regions occur outside shear-dominated patches where the thermal plumes merge successively with growing distances from the walls. As one approaches the mid-plane of the convection cell, they tend to form the turbulent superstructure pattern of convection. We have determined that the area fraction of the incoherent regions remains nearly constant at approximately 60% of $A$ for the whole Rayleigh number range. The insensitivity of the volume fractions with respect to the Rayleigh number suggests that this skeleton of upwelling (downwelling) incoherent regions could be a relic from the weakly nonlinear regime of convection at much lower Rayleigh numbers, which itself arises from the onset of convection by a linear primary instability, filling the whole domain with convection rolls.
We have varied the threshold for this analysis from $0.5U^{h}_{\rm rms}(\delta_{T})$ to $2U^{h}_{\rm rms}(\delta_{T})$. While the incoherent fractions do depend on the threshold when its variations are large, they are practically independent of the Rayleigh number even for such large variations stated above. This supports our choice of $U^{h}_{\rm rms}(\delta_{T})$ as a physically meaningful threshold.
@@
Figure 8: Boundary layer flow structure for $Ra=10^{10}$ at the top and bottom walls. (a) Decomposition of a snapshot into coherent and incoherent (in gray) flow regions at $z=\delta_{T}$. For the coherent shear-dominated regions, we overlay horizontal streamlines and indicate their local flow direction. (b) Corresponding temperature field in plane at $z=0.1\approx 25\delta_{T}$. Gray area corresponds to $T>T_{0}=0.5$. (c) Same snapshot with the decomposition similar to (b) for $z=H-\delta_{T}$ close to the top wall. (d) Corresponding temperature field at $z=0.9\approx 25\delta_{T}$. Gray area corresponds now to $T<T_{0}$. Note that all plots are shown at the coarse resolution of $100\times 100$ square boxes (which were used to calculate the local mean flow magnitude and orientation) and not at the original spectral resolution of the DNS run.
We can now return to the fluctuation analysis which is conditioned on coherent and incoherent regions in the following. Figure 9 replots the rms profiles of full and horizontal velocity and temperature profiles for Rayleigh numbers $Ra=10^{8}$, $10^{9}$, and $10^{10}$. We have chosen these three Rayleigh numbers of our series to provide a one-to-one comparison with DNS data in a closed cylindrical cell at $\Gamma=1$ of Scheel & Schumacher (2017). They are also shown in the figure. Vertical profiles, which have been taken over the full cross section (denoted as case G4 in the following), are shown in the left column of figure 9. Profiles conditioned on shear-dominated regions are displayed in the middle column (case G4C), while those for the cylindrical cross section of the closed container (case G1) are shown in the right column. From the bottom row of the figure, it is clear that the temperature profiles of G4, G4C, and G1 for all three $Ra$ agree. This suggests that the temperature boundary layers are alike in all cases. This is different for the velocity field, for which the horizontal velocity fluctuations (displayed in top row) show a clear trend. The thickness scale decreases from G4 to G4C and even more from G4C to G1. The close agreement of G4C and G1 clearly supports the dominance of shearing motion in the boundary regions in closed cylindrical cells, imposed by the prominent large-scale circulation. It is in line with a reduced fluctuation thickness. For fluctuations with respect to the full velocity field, we do detect a decrease of the thickness from G4 to G4C, but not from G4C to G1 for the two lower $Ra$. We suspect that this might be caused by prominent coherent up- and downwelling motions at the side walls for the lower $Ra$ which effectively enhance the thickness (Schumacher & Scheel, 2016).
Unlike G1 and G4C, for which the velocity and temperature boundary layers have comparable thicknesses, the G4 case shows that the velocity boundary layer is much thicker than the temperature thickness, suggesting a different mechanism in G4. We recall that the notion of the velocity boundary layer is only nominal in the sense that they are based on fluctuation profiles and the mean velocity variation within that region is quite small (see figure 4).
Figure 9: Vertical profiles of the velocity and temperature fluctuations for Rayleigh numbers $Ra=10^{8}$, $10^{9}$, and $10^{10}$. The figure compares averages over the full cross section at $\Gamma=4$ (left column) with averages over the coherent shear-dominated regions (middle column) and with averages over the full circular cross section in closed cylindrical cell at $\Gamma=1$ (right column) from Scheel & Schumacher (2017). Panels (a,b,c) are for $U^{h}_{\rm rms}$, panels (d,e,f) for $U_{\rm rms}$, and panels (g,h,i) for $T_{\rm rms}$. The dashed lines in each panel mark the location of the first local maximum away from the wall, and indicate the corresponding fluctuation thickness. The color coding, which is indicated in panel (c), holds for all panels.
Figure 10: Boundary layer thickness-based Rayleigh number $Ra_{\delta}$ versus Rayleigh number $Ra$. We have collected for comparison data from other simulations including Scheel & Schumacher (2017) and Iyer et al. (2020); also plotted are experimental data from Fleischer & Goldstein (2002) and Niemela et al. (2000). Note that all listed data from previous sources use $\delta_{T}=1/(2Nu)$ to calculate the boundary layer thickness-based Rayleigh number $Ra_{\delta}$. Only the present data use $\delta_{T,{\rm rms}}$ for the determination of $Ra_{\delta}$. The corresponding Rayleigh numbers $Ra_{\delta}=Ra_{\delta,{\rm rms}}$ are found in table 1.
7 Final discussion